What is the mass, in grams, of each of the following?

2.0×10−3 mol Sb

3.51×10−2 mol Ba

44.1 mol Xe

1.1 mol W
Answer #1

we know

mole = mass / molar mass

mass = moles * molar mass

1. mass of Sb = 2*10-3 mol * 121.76 g/mole = 0.243 g

2. mass of Ba = 3.51 * 10-2 mol * 137.33 g/mole = 4.82 g

3. mass of Xe = 44.1 mol * 131.29 g/mole = 5789.889 g

4. mass of W = 1.1 mol * 183.8 g/mole = 202.18 g